# SFFT-based Homogenization: Using Tensor Trains to Enhance FFT-Based Homogenization

### v1

@@ 4.2 The Zip-Up Algorithm
The ”Zip-Up” algorithm can be used to convert any quantum circuit into a TTO through a combination of tensor contractions and SVDs [33]. In the following, we will explain the algorithm by transferring the reduced QFT into the SFFT.
Before applying the procedure on the QFT quantum circuit, we must perform preliminary preparations by decomposing the controlled quantum gates and recombining them in a slightly more favourable manner. In this case, this means separating the controlled Phase-Gates into a ’control’ and a ’phase’ components. Mathematically, this is achieved by representing the full gate as a separated inner product of the following form
@@
This results in a combination of a Copy Gate and a 3-legged Phase Gate (or simply 3-phase gate), as shown in Figure 3 a). We then separate all controlled Phase gates into such pairs. The resulting copy gates enforce that all legs correspond to the same state. Therefore, the copy gates can be interchanged as long as the total incoming and outgoing dimensions, or ’legs’, remain fixed. Intuitively this can be described through a diagrammatic representation, see Figure 3 b).
These reordered copy gates can be recombined with the 3-phase gates, as shown in Figure 3 c), resulting in a novel 4-legged phase gate (or 4-phase gate). It should be noted that although we recombine the copy and 3-phase gates, the incoming and outgoing dimensions have changed.
With our prepared mathematical construct, we can reformulate the reduced QFT into QTT format using the Zip-Up algorithm. Figure 2 a) shows the diagrammatic representation of the Zip-Up algorithm, while the specific transformations required to perform the block contractions are shown in b).
The algorithm begins by contracting the tensor on the least significant qubit with its right neighbour along their shared index. In the second step, this new tensor is reshaped into a matrix. The incoming and outgoing dimensions of the matrix are determined by specific leg combinations. Since the algorithm starts from the bottom and goes to the top, the vertical as well as the downward legs get combined into the outgoing dimension, while the upward legs form the incoming dimension. As a next step, a SVD is performed on this matrix. The $V$-matrix remains, while the $U$ and $S$ matrices get recombined. As a last sub-step, the remaining two matrices get reshaped back into tensor form, where special care needs to be taken into separating the legs into their original form.
Figure 2: a) The first few steps of Zip-Up Algorithm in example of the QFT for the full Quantum Circuit; b) The detailed transformations needed within one step of the Zip-Up algorithm.
In the next iteration of the procedure, the resulting $V$-tensor remains at the original position, while the $US$-tensor gets contracted with the two tensors acting on the next upper neighbouring qubit, see Figure 2. This procedure is repeated until the most significant qubit is reached. Now the procedure restarts by recombining the $V$-tensor acting on the least significant qubit with its right neighbour ($V_{4}$ in Figure 2). The full down-top procedure is performed until there is only one tensor acting per qubit left.
@@ 4.3 SFFT-based Homogenization
To derive the SFFT-based Homogenization algorithm, the corresponding states and operators of the FFT-based algorithm 1 must be transferred into QTT or QTTO format. The only naturally occurring operators are the FFT and its inverse, which are simply replaced by the novel (inverse) SFFT operator, denoted as $\mathcal{S}^{(-1)}$. The other states - namely, the local strain $\varepsilon(\mathbf{x})$, the mean strain $\mathbf{E}$, the Green-Eshelby Operator $\mathnormal{\hat{\Gamma}}_{0}(\mathbf{q})$ and the local and isotropic reference stiffness $\mathbf{C(x)}$ and $\mathbf{C}_{0}$ must be formulated in QTT format. To achieve this, we begin by discretizing the $\mathbf{x}$ or $\mathbf{q}$ dependent states in a binary fashion. This means that, instead of discretizing a function $t(x)$using a single grid variable $k$, according to
@@
Although the function is evaluated at the same grid points, the resulting tensor has a different number of physical indices and corresponding dimensions. The binary discretization method yields an $n$-order tensor with dimensionality $d=2$ per index, which mirrors the physical indices required for the QTT format. Once the high-order tensor is obtained, the QTT cores can be separated easily by applying the (randomized) TT-SVD or Tensor-Cross-Approximation [29, 34], as detailed in Appendix D.
In our multidimensional case, the dimensions are processed sequentially, resulting in QTT representations where the cores are grouped by dimension. We assume $N=2^{n}$ discretization points per dimension, corresponding to $n$ QTT cores (similar to $n$ qubits) per dimension.
Two constant operators are required for the algorithm’s iterative procedure: the mean-strain $\mathbf{E}$ and the isotropic reference stiffness $\mathbf{C}_{0}$. These special cases can easily be formulated by a QTT of rank one by directly generating the cores using $(2\times 2)$ identity matrices, resulting in core shapes of $(2\times 1\times 2\times 1)$. Multiplying one of the cores with the constant value associated with the respective functions results in their respective QTT.
Until now, we have only discussed the internal structures of the tensors in the context of the required discretization. However, the tensors in question generally yield multidimensional outputs, introducing additional physical indices. Here, we are considering the original tensors in voigt notation resulting in a simplified operator structure. One approach to handle these indices is to treat the states as vectors (or matrices), with their elements in QTT format. Although this approach is justifiable, exploiting the computational benefits of the TT optimization algorithms would not be feasible.
To obtain a full TT representation, we make a slight deviate from the QTT format by adding a single additional core at the beginning of the QTT. This core is permitted to deviate from the imposed dimensionality dictated by the QTT and serves as a control for the different components of the Voigt formulation. Thus, it holds the information about the Voigt arithmetic rather than the discretization of the function. Therefore, we refer to this version as the controlled QTT (cQTT). This approach generalizes naturally to incorporate controlled QTTOs (cQTTOs).
@@
Since the Green-Eshelby operator $\mathnormal{\hat{\Gamma}}_{0}$ resides in the momentum space, it must be treated differently. This is because of the required momentum-space convention of using the SFFT over the FFT. As discussed in section 4.1, a reordering of the cores is necessary. This involves two steps: 1) The indices per geometrical dimension $D$ of the initial discretization tensor, obtained through equation (13), must be reversed. 2) The tensor must be complex conjugated. Performing these additional steps for operators in the momentum space ensures the correct convention according to the SFFT, yielding our final form of the SFFT-based Homogenization algorithm 2.
@@ 5 Experiments
Figure 9: Maximal Strain Rank $r_{s}$ during the evaluation of the algorithm in dependence of the discretization order $n$ in a) 2D and b) 3D space.
We evaluated the performance of the SFFT-based Homogenization algorithm across various geometries to assess its applicability in diverse scenarios. The primary geometries used in this study are shown in Figure 4. Additional tests were also conducted on various geometries in both 2D and 3D space, as detailed in Appendix E.”
@@ 5.1 Setup and Preprocessing
Before executing the algorithm, all states and operators were converted into the cQTT format. The corresponding preprocessing time can be reduced by storing operators in lookup tables for future runs, as most of them can be reused, minimizing overhead. A more detailed exploration of preprocessing scalability and potential improvements can be found in Appendix D.
The geometries shown in Figure 4 were analysed in 2D, with two possible generalizations considered for 3D space. The first generalization extends the geometries by adding a trivial, non-varying z-component, turning the 2D Laminate into a 3D Laminate, the Square into a Pillar, and the Circle into a Spherical Pillar. The second generalization assumes that the geometries in Figure 4 are viewed identically from every direction, resulting in the Square becoming a Box and the Circle becoming a Sphere. The Laminate is limited to only a trivial generalization.
@@
$\Delta\bar{\sigma}=\mathop{\mathbb{E}}\left[\frac{|\sigma^{\mathrm{SFFT}}(\mathbf{x})-\sigma^{FFT}(\mathbf{x})|}{|\sigma^{FFT}(\mathbf{x})|}\right],$ (14)
where $\sigma^{\mathrm{SFFT}}$ and $\sigma^{FFT}$ are the stress fields obtained after convergence of the SFFT- and FFT-based Homogenization algorithms, respectively.
Unless otherwise specified, all experiments used a cQTT and cQTTO accuracy parameter of $\delta_{\mathrm{acc}}=10^{-6}$ and a convergence threshold of $\delta_{\mathrm{tol}}=10^{-4}$. The discretization followed a ’staggered grid’ approach, as outlined in Appendix F. For clarity, the discretization order $n$ is used throughout, where $N=2^{n}$ represents the number of discretization points per dimension $D$.
@@ 5.2 Performance Comparison
Figure 10: a) Stress field $\sigma_{yy}$ under tensile load and b) $\sigma_{xy}$ under shear load; c) relative error $\Delta\sigma_{yy}$ for the Square geometry and d) $\Delta\sigma_{xy}$ for the Circle geometry, evaluated with $n=8$ using the cQTT- and FFT-based approaches.
@@
The memory consumption results shown in Figure 5 highlight the exponential scaling of the FFT-based approach, in contrast to the geometry-specific advantages of the SFFT method in both 2D and 3D. For 1D-like geometries, such as the Laminate, the SFFT method significantly reduces memory usage. The Square Pillar and Box geometries demonstrate sub-exponential scaling, reflecting the adaptability of the SFFT method. In contrast, the Circle and Circular Pillar exhibit exponential scaling similar to the FFT method, while the Sphere performs even slightly worse than the FFT method. Notably, trivially generalized structures, such as the Pillar and Circular Pillar, show improved memory efficiency over their FFT baseline when compared to their 2D counterparts. This is a direct consequence of the additional dimension not increasing complexity for the SFFT due to its geometry-tailored approach.
@@
The iteration times, shown in Figure 6, reveal that for smaller discretization orders $n$, the FFT-based method is faster. However, the SFFT outperforms the FFT beyond a geometry-dependent crossover point. In 2D, the Laminate crosses over at $n=11$, followed by the Square at $n=15$. In 3D, the Laminate reaches the crossover substantially quicker at $n=7$, while both the Pillar and Circular Pillar geometries reach the crossover at a reasonable order of $n=10$. Furthermore, the plot indicates that for even higher discretization orders, the Box will exhibit a crossover around $n=15$, aligning with the 2D results. Although no crossover is expected for the Circle and Sphere, the Circular Pillar may experience one at very high discretizations, suggesting that trivially generalized 3D structures may always have a crossover point. While this discretization order would be far to high for any real-world application, this may suggests that trivially generalized 3D structures may always exhibit a crossover point due to the negligible third dimension. It should be noted that the FFT baseline benefits from highly optimized libraries, while similarly optimized code is not yet available for the SFFT-based Homogenization. Thus, we expect that further optimizations to the SFFT implementation will noticeably reduce the order at which crossover points occure.
@@
Furthermore, it is worth noting that the iteration times between the Square and Circle and their trivial extensions, the Pillar and Circular Pillar, respectively, are almost negligible. The minor difference is primarily due to the non-negligible contribution of the Green-Eshelby Operator in 3D.
@@
Figure 7 shows the error dependence on the discretization order $n$. For the Laminate, the analytic solution is achieved after just one iteration, resulting in a significantly reduced relative error. For other geometries, the relative error stabilizes around $10^{-5}$ for both 2D and 3D cases, well within the acceptable range for industrial applications. The error remains close to the rounding accuracy of the cQTTs of $\delta_{\mathrm{acc}}$, see equation (12).
@@
Figure 8 shows that the SFFT and ISFFT cQTTO ranks converge to $r_{\mathrm{G}}=10$ after $n=8$ in both 2D and 3D, indicating the efficiency and scalability of the respective operators. The Green-Eshelby operator exhibits approximately linear rank growth across discretization and dimension, further demonstrating the favourable scaling of the non-geometry-dependent parts of the SFFT approach. Therefore, the applicability of SFFT-based Homogenization is limited by the compressibility of the underlying geometry. The Circle and its 3D counterparts consistently exhibit exponential rank growth, highlighting the inherent incompatibility of there geometries with the SFFT approach. This is further supported by Figure 9, where similar growth patterns are observed for the strain rank $r_{\mathrm{S}}$ of the respective geometries.
@@
These results suggest that it may be possible to quickly estimate the applicability of the SFFT-based Homogenization by compressing the geometries at two different discretization orders. This approach could provide an initial indication of both the rank and its growth behaviour, helping to assess whether a geometry is suitable for our approach.
@@
Finally, it is worth noting that in the FFT-based Homogenization, iteration times are primarily limited by the FFT computation itself. In contrast, our SFFT-based approach shifts the computational bottleneck to the matrix-vector product between the Green-Eshelby cQTTO and the strain cQTT. This operation typically accounts for more than 95% of the total computation time due to the relatively high ranks of both the Green-Eshelby operator and the strain during the iterative procedure. This distinction underscores a key area for potential optimization in our method. Since the full scalability limits of the SFFT have not yet been reached, improving the efficiency of this matrix-vector product could lead to substantial performance gains.

### v3

@@ 4.2 The Zip-Up Algorithm
Figure 2: a) The first steps of the Zip-Up algorithm in example of the reduced QFT quantum circuit; b) The detailed transformations needed within one step of the Zip-Up algorithm.
The ’Zip-Up’ algorithm can be used to convert any quantum circuit into a TTO through a combination of tensor contractions and SVDs [44], as shown in Figure 2. In the following, we will explain the algorithm by transferring the reduced QFT into the SFFT.
Before applying the procedure on the QFT quantum circuit, we must perform preliminary preparations by decomposing the controlled quantum gates and recombining them in a slightly more favourable manner. In this case, this means separating the controlled Phase-Gates into a ’control’ and a ’phase’ component. Mathematically, this is achieved by representing the full gate as a separated inner product of the following form
@@
This results in a combination of a Copy Gate and a 3-legged Phase Gate (or simply 3-Phase Gate), as shown in Figure 3 a). We then separate all controlled Phase Gates into such pairs. The resulting Copy Gates enforce that all its inputs and outputs correspond to the same state. Therefore, the Copy Gates can be interchanged as long as the total incoming and outgoing dimensions, or ’legs’, remain fixed. Intuitively, this can be illustrated through a diagrammatic representation, as visualized in Figure 3 b).
These reordered Copy Gates can be recombined with the original 3-Phase Gates, resulting in novel 4-Phase Gates, as shown in Figure 3 c). It should be noted that although we recombine the Copy and 3-Phase Gates, the incoming and outgoing dimensions have changed. With our prepared construct, we can reformulate the reduced QFT in TTO format using the Zip-Up algorithm.
The algorithm, depicted in Figure 2 a), begins by contracting the tensor of the least significant qubit with its right neighbour along their shared index. In the second step, this new tensor is reshaped into a matrix, as shown in Figure 2 b). The incoming and outgoing dimensions of the matrix are determined by the specific leg combinations. Since the algorithm starts from the bottom and goes to the top, the vertical as well as the downward legs get combined into the outgoing dimension, while the upward legs form the incoming dimension. As a next step, a SVD is performed on this matrix. The $V$ matrix remains, while the $U$ and $S$ matrices get recombined into a new $US$ matrix. As the final sub-step, the remaining two matrices are reshaped back into tensor form, with special care taken to separate the legs back into their original configuration.
In the next iteration of the procedure, the resulting $V$-tensor remains at the original position, while the $US$-tensor gets contracted with the two tensors acting on the next upper neighbouring qubit. This procedure is repeated until the most significant qubit is reached. Now the procedure restarts by recombining the $V$-tensor acting on the least significant qubit with its right neighbour ($V_{4}$ in Figure 2). This down-top process continues until only one tensor acting per qubit is left.
@@ 4.3 SFFT-based Homogenization
To derive the SFFT-based Homogenization algorithm, the corresponding operands and operators of the FFT-based algorithm 1 must be transferred into TT and TTO format, respectively. The only naturally occurring operators are the FFT and its inverse. They are simply replaced by the novel SFFT operator $\mathcal{S}$ as well as its inverse ISFFT operator $\mathcal{S}^{-1}$. The operands - namely, the local strain $\varepsilon(\mathbf{x})$, the mean strain $\mathbf{E}$, the Green-Eshelby operator $\mathnormal{\hat{\Gamma}}_{0}(\mathbf{q})$ and the local and isotropic reference stiffnesses $\mathbf{C(x)}$ and $\mathbf{C}_{0}$ must be formulated in TT format. To achieve this, we begin by discretizing the $\mathbf{x}$ or $\mathbf{q}$ dependent operands in a binary fashion. This means that, instead of discretizing a function $t(x)$ using a single grid variable $k$, according to
@@
Although the function is evaluated at the same grid points, the resulting tensor has a different number of physical indices and corresponding dimensions. The binary discretization method yields an $n$-order tensor with dimensionality $d=2$ per index, which mirrors the physical indices required for our used TT format. Once the high-order tensor is obtained, the TT cores can readily be separated by successively applying the (randomized) SVD algorithm or the Tensor-Cross-Approximation algorithm [37, 45], as detailed in Appendix D.
In our multidimensional case, dimensions are processed sequentially, grouping cores by dimension in its TT representations. We assume $N=2^{n}$ discretization points per dimension $D$, corresponding to $n$ cores (similar to $n$ qubits) per dimension.
Two constant operands are required for the algorithm’s iterative procedure: the mean-strain $\mathbf{E}$ and the isotropic reference stiffness $\mathbf{C}_{0}$. These special cases can be easily represented by a TT of rank one by first generating $(2\times 2)$ identity matrices and reshaping them into the appropriate core shapes of $(2\times 1\times 2\times 1)$. Multiplying one of the cores by the constant value of the associated operand results in their respective TT.
Until now, we have only discussed the internal structures of the tensors in the context of the required discretization. However, the tensors in question generally yield multidimensional outputs, introducing additional physical indices. Here, we are considering the original tensors in Voigt notation resulting in a simplified operator structure. One approach to handle these indices is to treat the operands as vectors (or matrices), with their elements in TT format. Although this approach is justifiable, exploiting the computational benefits of the TT optimization algorithms would not be feasible. To obtain a full TT representation, we make a slight deviate from the currently used TT format by adding a single additional core at the beginning of the TT. This core is permitted to deviate from the imposed dimensionality of $d=2$ and serves as a control for the different components of the Voigt formulation. Thus, it holds the information about the Voigt arithmetics rather than the discretization of the function. Therefore, we refer to this version as the controlled Quantized TT. This approach generalizes naturally to incorporate controlled Quantized TTOs. However, for the sake of readability, we will not introduce a new abbreviation and will continue to use TT(O) for this specialization.
@@
Further, since the Green-Eshelby operator $\mathnormal{\hat{\Gamma}}_{0}(\mathbf{q})$ resides in the momentum space, it must be treated separately. This is due to the momentum space convention required to use the reduced QFT as the underlying operator for the SFFT. As discussed in Section 4.1, a reordering of the cores is necessary. This involves two steps: 1) The indices per geometrical dimension $D$ of the initial discretization tensor, obtained through equation (13), must be reversed. 2) The tensor must be complex conjugated. Performing these additional steps for operators in the momentum space ensures the correct convention according to the SFFT, yielding our final form of the SFFT-based Homogenization algorithm 2.
Examining Table 2, it becomes clear that the efficiency of the SFFT-based algorithm is highly dependent on the rank dimensions of the operators involved and the strain field. The total time complexity is given by
$\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}\left(r_{\mathrm{C}}^{3}+r_{\varGamma}\right)\Big{\}}.$ (14)
with the rank $r_{\mathrm{S}}$ of the strain TT $\varepsilon_{m}$, the rank $r_{\mathrm{C}}$ of the reduced stiffness TTO $\left[\mathbf{C-C}(\mathbf{x})\right]$ and the rank $r_{\varGamma}$ of the Green-Eshelby TTO $\hat{\varGamma}_{0}$. Due to brevity, we will only state the time complexity of the SFFT-based Homogenization algorithm and provide a detailed derivation in Appendix E. Our results show that as long as the contributing ranks remain manageable, an exponential speedup is achieved in both the number of discretization points $N=2^{n}$ and geometric dimensionality $D$, in comparison to the FFT-bottlenecked Homogenization method. The geometry dependent term within the bracket of relation (19) can be computed during the offline phase, allowing for an informed estimate on the approaches applicability before running the full algorithm.
@@ 5 Experiments
Figure 9: Maximal Strain Rank $r_{\mathrm{S}}$ during the evaluation of the algorithm in dependence of the discretization order $n$ in a) 2D and b) 3D space.
Figure 10: a) Stress field $\sigma_{yy}$ under tensile load, and b) point-wise relative error $\Delta\sigma_{yy}$ for the Square geometry. c) Stress field $\sigma_{xy}$ under shear load, and d) point-wise relative error $\Delta\sigma_{xy}$ for the Square geometry. The evaluation used $n=8$ for the SFFT- and FFT-based algorithms.
We evaluated the performance of the SFFT-based Homogenization algorithm across various geometries to assess its applicability in diverse scenarios. The primary geometries used in this study are shown in Figure 4. Additional tests were conducted on various geometries in both 2D and 3D space, as detailed in Appendix F.
@@ 5.1 Setup and Preprocessing
Before executing the algorithm, all operands and operators were converted into the TT format. The corresponding preprocessing time can be reduced by storing operators in lookup tables for future runs, as most of them can be reused, minimizing overhead. A more detailed exploration of preprocessing scalability and potential improvements can be found in Appendix D.
The geometries shown in Figure 4 were analysed in 2D, with two generalizations considered for 3D. The first generalization extends the geometries by adding a trivial, non-varying z-component, turning the 2D Laminate into a 3D Laminate, the Square into a Pillar, and the Circle into a Circular Pillar. The second generalization assumes that the geometries in Figure 4 are viewed identically from every direction, resulting in the Square becoming a Box and the Circle becoming a Sphere. The Laminate is limited to a trivial generalization only.
@@
$\Delta\bar{\sigma}=\mathop{\mathbb{E}}\left[\frac{|\sigma^{\mathrm{SFFT}}(\mathbf{x})-\sigma^{\mathrm{FFT}}(\mathbf{x})|}{|\sigma^{\mathrm{FFT}}(\mathbf{x})|}\right],$
where $\sigma^{\mathrm{SFFT}}$ and $\sigma^{\mathrm{FFT}}$ are the stress fields obtained after convergence of the SFFT- and FFT-based Homogenization algorithms, respectively.
Unless otherwise specified, all experiments used a TT and TTO accuracy parameter of $\delta_{\mathrm{acc}}=10^{-6}$ and a convergence threshold of $\delta_{\mathrm{tol}}=10^{-4}$. The discretization followed a ’staggered grid’ approach, as outlined in Appendix G. For clarity, the discretization order $n$ is used throughout, where $N=2^{n}$ represents the number of discretization points per dimension $D$.
@@ 5.2 Performance Comparison
The memory consumption results, shown in Figure 5, highlight the exponential scaling of the FFT-based approach, compared to the geometry-specific SFFT method in both 2D and 3D. For 1D-like geometries, such as the Laminate, the SFFT method significantly reduces memory usage. The Square, Pillar and Box geometries demonstrate subexponential scaling, reflecting the adaptability of the SFFT method. In contrast, the Circle and Circular Pillar exhibit exponential scaling similar to the FFT method, while the Sphere performs even slightly worse than the FFT-based approach. Notably, trivially generalized structures, such as the Pillar and Circular Pillar, show similar scaling as their 2D counterpart. This is a direct consequence of the additional (trivial) dimension not notably increasing the runtime for the SFFT-based method due to its geometry-tailored approach.
@@
The iteration times, shown in Figure 6, reveal that for smaller discretization orders $n$, the FFT-based method is faster. However, our approach outperforms the FFT-based algorithm beyond a geometry-dependent crossover point. In 2D, the Laminate crosses over at $n=11$, followed by the Square at $n=15$. In 3D, the Laminate reaches the crossover substantially quicker at $n=7$, while both the Pillar and Circular Pillar geometries reach the crossover at a reasonable order of $n=10$. Furthermore, the plot indicates that for even higher discretization orders, the Box will exhibit a crossover around $n=15$, aligning with the 2D results. Although no crossover is expected for the Circle and Sphere, the Circular Pillar may experience this point at very high discretizations. While the corresponding discretization order would be far too high for any real-world application, this may indicate that trivially generalized 3D structures consistently display a crossover point because of their negligible third dimension. Finally, it should be noted that the FFT baseline benefits from highly optimized libraries, while similarly optimized code is not yet available for the SFFT-based Homogenization algorithm. Thus, we anticipate that further optimizations to our implementation will significantly reduce the discretization order at which crossover points occur.
@@
Furthermore, it is worth noting that the iteration times between the Square and Circle and their trivial extensions - the Pillar and Circular Pillar - respectively, are almost negligible. The minor difference is primarily due to the non-negligible contribution of the Green-Eshelby operator in 3D. This mirrors the behaviour observed in the memory scaling experiments.
@@
The relative error $\Delta\overline{\sigma}$ in dependence on the discretization order $n$, is shown in Figure 7. For the Laminate, the analytic solution is achieved after just one iteration, resulting in a significantly reduced error. For all other geometries, the relative error stabilizes around $10^{-5}$ for both 2D and 3D cases, well within the acceptable range for industrial applications. Noticeably, the error remains close to the accuracy of the TT rounding procedure of $\delta_{\mathrm{acc}}=10^{-6}$, see equation (12).
@@
Figure 8 shows that the SFFT and ISFFT TTO ranks converge to $r_{\mathrm{G}}=10$ after $n=8$ in both 2D and 3D, indicating the efficiency and scalability of the respective operators. The Green-Eshelby operator exhibits approximately linear rank growth across discretization and dimension, further demonstrating the favourable scaling of the nongeometry-dependent parts of the SFFT approach. Therefore, the applicability of SFFT-based Homogenization is limited by the compressibility of the underlying geometry. The Circle and its 3D counterparts consistently exhibit exponential rank growth, highlighting the inherent incompatibility of these geometries with the current SFFT approach. This is further supported by Figure 9, where similar growth patterns are observed for the strain rank $r_{\mathrm{S}}$ of the respective geometries. In these exponential scaling cases, the geometry-dependent term of the time complexity relation (19) dominates due to the considerable increase of $r_{\mathrm{C}}$. This leads to $\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}r_{\mathrm{C}}^{3}\Big{\}}$, where both ranks introduce a cubic contribution, resulting in a rapidly growing compute time. Conversely, in subexponential scaling cases, the geometry dependent term is typically negligibly, resulting in an approximate time complexity of $\mathcal{O}\Big{\{}nD\,r_{\mathrm{S}}^{3}r_{\varGamma}\Big{\}}$, which is primarily governed by the Green-Eshelby TTO and exhibits improved rank scaling behaviour.
@@
Moreover, these results suggest that it may be possible to quickly estimate the applicability of our method by compressing the geometries at two different discretization orders. This approach could provide an indication of the rank and its growth behaviour, helping to assess whether our method is suitable for a given geometry.
@@
Finally, it should be noted that in the FFT-based algorithm, iteration times are primarily limited by the FFT computation itself. In contrast, our SFFT-based approach shifts the computational bottleneck to the matrix-vector product between the Green-Eshelby TTO and the strain TT. This operation typically accounts for over 95% of the total computation time, due to the high ranks of both the TT and TTO involved. This highlights a key area for potential optimization. As the full scalability limits of the SFFT remain unexplored, enhancing the efficiency of this matrix-vector product could yield significant performance gains.